There is a second part to this question asking how far does it fall in the first 1.9 seconds of its flight??
Help Required information Problem 3.00E1 DEPENDENT MULTI-PART PROBLEM-ASSIGN ALL PARTS NOTE: This is a multi-part question. Once an answer is submitted, you will be unable to return to this part A steel ball is dropped from a diving platform (with an initial velocity of zero). Use the approximate value of g as 10 m/s? to solve the following problem 47 Problem 3.00E1.a-Calculating the velocity of a dropped object What is the velocity of the ball 0.5 seconds after its release? (You must provide an answer before moving to the next part) The velocity of the ball is m/s.
Answer #1

initial velocity u =0

final velocity v = u+gt

t= 0.5 s and g = 10 m/s2

final velocity after 0.5 s v = 0 + 10*0.5 = 5 m/s